The number of requests for assistance received by a towing service is a Poisson process with rate ? = 4 per hour.

a) If the operators take a 30 min break for lunch, what is the probability that they do not miss any calls for assistance?

b) Assuming they work 10 total hours on a particular day. What is the probability that assist less than 30 people (consider and approximation to help you solve this part)?

c) Calculate the average waiting time between requests (consider a continuous distribution).
Answer #1

a) Here, ? = 2 in 30-mins. The operators will not miss any calls if none came in. Thus, the required probability is

P(X = 0) = e-22 e-יגג 지 = 0, 135335

b) Here, ? = 40 in 10-hours. Therefore,

P(X <30) P(Xs29) 29 e 4040 = 0043228682

c) Mu = 1/4 = 0.25